OpenStudy (anonymous):

The area of a given hexagon is equal to the area of an equilateral triangle whose perimeter is 36 inches. Find the length of a side of the regular hexagon.

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~\to~OpenStudy!!}\hspace{-310pt}\color{black}{Welcome~\to~OpenStudy!!}\hspace{-307.1pt}\color{lime}{Welcome~\to~\color{lime}{Open}}\color{lime}{Study!!}\)

OpenStudy (yanasidlinskiy):

area of the triangle = 1/2 * 12 * sqrt(12^2 - 6^2) area of regular hexagon = 3*sqrt3 * s^2 ------ 2 where s = length of the side equate these and solve for s